Find the sum: 8/15 + (-2/3)
Find the sum: -3 1/4 + (-2 2/5)
Find the difference: 5/8 - 3/4
Find the difference: -4/10 - (-1/6)
Find the difference: -6 2/3 - 4 3/5

Any help will be appreciated! Thanks!

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Find the sum: 8/15 + (-2/3) = (8/15)-(10/15) = -2/15
---------------------------------------------------------
Find the sum: -3 1/4 + (-2 2/5) = (-13/4)-(12/5)
= (-65/20)-(48/20) = -113/20
---------------------------------------------------------
Find the difference: 5/8 - 3/4 = (5/8)-(6/8) = -1/8
---------------------------------------------------------
Find the difference: -4/10 - (-1/6) = (-12/30)+(5/30) = -7/30
---------------------------------------------------------
Find the difference: -6 2/3 - 4 3/5 = (-20/3)-(23/5)
= (-100/15)-(69/15)
= -169/15
==========================================================
Cheers,
Stan H.

Answer by checkley77(12844)

Find the sum: 8/15 + (-2/3)
8/15-10/15=-2/15 ans.
Find the sum: -3 1/4 + (-2 2/5)
(-12+1)/4-(2*5+2)/5
-11/4-(10+2)/5
-11/4-12/5
(-11*5-12*4)/4*5
(-55-48)/20
-103/20 ans.
Find the difference: 5/8 - 3/4
5/8-6/8=-1/8 ans.
Find the difference: -4/10 - (-1/6)
-4/10+1/6
(-4*5+1*3)/30
(-20+3)/30
-17/30 ans.
Find the difference: -6 2/3 - 4 3/5
(-6*3+2)/3-(4*5+3)/5
(-18+2)/3-(20+3)/5
-16/3-23/5
(-16*5-23*3)/(3*5)
(-80-69)/15
149/15 ans.